1. A ball of mass 0.6 kg is dropped from a height of 1.8m onto a solid floor. Each time the ball
bounces on the floor it loses 10% of its speed.
a) Work out how much momentum was absorbed by the floor in the first bounce?
b) Show that the ball first fails to reach a height of 1 m after the third bounce.
1
Expert's answer
2021-09-08T16:14:30-0400

"g\approx 10\frac{m}{s^2}"

"m= 0.6kg"

"h_1= 1.8m"

"v_n-\text{ball speed when reaching ground level}"

"v_n'\text{-ball velocity after hitting the ground}"

"a)"

"\text{According to the law of conservation of energy:}"

"mgh_1= \frac{mv_1^2}{2}"

"v_1=\sqrt{2gh_1}= \sqrt{2*10*1.8}= 6\frac{m}{s}"

"v_1' = v_1-v_1*10 \%=0.9v_1=0.9*6=5.4\frac{m}{s}"

"p_1 = mv_1"

"p'_1=mv'_1"

"\Delta p = p'_1-p=m(v'_1-v)=0.6*(6-5.4)= -0.36kg\frac{m}{s}"

"\text{Answer: }-0.36kg*\frac{m}{s}"

"b)"

"v_1'=0.9v_1"

"v_2 = v'_1"

"v_2'=0.9v_2=0.9^2v_1"

"v_3 = v'_2"

"v_3'=0.9v_3=0.9^3v_1=0.9^3*6=4.374\frac{m}{s}"

"mgh_3 = \frac{mv'^2_3}{2}"

"h_3 = \frac{v'^2_3}{2g}=\frac{4.374^2}{2*10}\approx0.96m"

"\text{Answer: }h_3 = 0.96m<1m"